A motarboat can maintain a constant speed of 50 mph relative to the water. the boat makes a trip upstream to a certain point in 51 minutes; the return trip takes 49 minutes. what is the speed of the current?
Answer by mananth(16946)

upstream it takes 51 minutes
let the speed of current be x mph
speed of boat will be 50-x
speed * time = distance
=(50-x) * 51/60
downstream it takes 49 minutes
speed of boat = 50+x
Distance = (50+x)*49/60
Since the distance traveled both ways is the same
(50-x)*51/60 = (50+x)*49/60
(50-x)*51 = (50+x)*49
2550-51x=2450+49x
2550-2450=51x+49x
100=100x
x=1 mph which is the speed of the current
Ananth